OpenStudy (anonymous):

given the function f(x) = x^3+2x^2-3x-5 what is resulting function when f(x) is shifted to the right 1 unit

OpenStudy (anonymous):

@frogs29 @douglaswinslowcooper

OpenStudy (anonymous):

@ganibl

OpenStudy (anonymous):

Shifted to the right one unit. g(x) = f(x-1), substitute (x-1) for x in f(x) to get g(x). e.g., f(x) = m x + b f(0) = b g(x) = m (x-1) + b g(1) = b Old f values now occur for g(x) at x+1. Shifted one unit to the right.

OpenStudy (anonymous):

No idea whats happening here @douglaswinslowcooper

OpenStudy (anonymous):

Given equation f(x) = x^3+2x^2-3x-5 It is shifted by 1 unit which means we need to calculate f(x + 1)

OpenStudy (anonymous):

f(x + 1) = (x + 1)^3 + 2 (x + 1)^2 -3(x + 1) - 5 x^3 + 3x^2 + 3x + 1 + 2(x^2 + 2x + 1) -3x - 3 - 5 x^3 + 3x^2 + 3x + 2x^2 + 4x -3x -5 x^3 + 5x^2 +4x - 5

OpenStudy (anonymous):

The resultant equation will be x^3 + 5x^2 + 4x - 5 when f(x) is shifted by one unit to the right

OpenStudy (anonymous):

Oh wow.. Yes much better thank you!!! can youhelp with another please?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = x^4 + 3x^3 -5x^2 + 2x -2. From x=-1 to x=1

OpenStudy (anonymous):

@ganibl

OpenStudy (anonymous):

anything?